# Understanding the Skill Gap in Recurrent Language Models: The Role of the Gather-and-Aggregate Mechanism

### v1

@@ Abstract
State-space models (SSMs) have recently emerged as efficient recurrent alternatives to Transformers for processing long sequences with fixed state sizes. However, this efficiency comes at the cost of reduced capabilities in certain algorithmic tasks, such as retrieval of past context. In this work, we examine how such in-context retrieval operates within Transformer- and SSM-based language models. We find that both architectures develop the same fundamental Gather-and-Aggregate (G&A) mechanism. A Gather Head first identifies and extracts relevant information from the context, which an Aggregate Head then integrates into a final representation. Across both model types, G&A concentrates in just a few heads, making them critical bottlenecks even for benchmarks that require a basic form of retrieval. For example, we show that disabling a single Gather or Aggregate Head of a pruned Llama-3.1-8B severely degrades its ability to retrieve the correct answer letter in MMLU, reducing its accuracy from 66% to 25% (random guessing). Moreover, this finding suggests that in-context retrieval can obscure the limited knowledge demands of certain tasks. Although the pruned model performs well on MMLU with functional retrieval, it struggles on other knowledge benchmarks. This G&A bottleneck extends to GSM8K, BBH, and dialogue tasks—all areas where SSMs typically underperform. Given the significance of G&A capability in performance, we show that the known retrieval challenges in SSMs manifest in how they implement G&A within these heads, leading to smoother attention patterns rather than the sharp token transitions that effective G&A relies on. Thus, while a gap exists between Transformers and SSMs in implementing in-context retrieval, it is limited to just a few heads, rather than a property of the entire language model. This insight points to a possible unified explanation for empirical performance differences between Transformers and SSMs while also highlighting ways to effectively combine their strengths. For example, we show that in pretrained hybrid models, where SSMs are interspersed with a few attention layers, the attention components naturally take on the role of Aggregate Heads. Similarly, in a pretrained pure SSM, replacing a single G&A head with an attention-based variant significantly improves retrieval capabilities, leading to higher benchmark scores.
1 Introduction
The rise of Transformers has reshaped modern language modeling, driving breakthroughs in natural language processing. Yet, Transformers’ quadratic scaling with sequence length has renewed interest in recurrent alternatives, particularly state-space models (SSMs). SSMs offer linear scaling with a fixed-size memory while remaining competitive across many tasks. However, replacing a large, explicit history cache with a compact state representation poses a fundamental tradeoff: SSM-based language models often struggle with the skill of “in-context retrieval”—the ability to precisely locate and extract tokens presented earlier in the sequence, such as recalling a phone number mentioned several paragraphs ago (Jelassi et al., 2024). This “algorithmic” challenge, which Transformers handle more effectively through direct attention to past tokens, has been identified as a key factor in the performance gap between Transformer models and SSM variants (Arora et al., 2023).
While previous work has explored retrieval limitations in SSMs (Wen et al., 2024; Jelassi et al., 2024; Waleffe et al., 2024), these studies have either treated language models as black boxes or focused on simplified theoretical setups. Meanwhile, research on Transformers has shown that algorithmic capabilities, including retrieval, are often concentrated in specific attention heads (Lieberum et al., 2023; Wu et al., 2024). This raises a key question: Can the performance gap between Transformer- and SSM-based language models be traced to a few algorithmic capabilities within a small subset of heads? Clarifying this will enable a more targeted comparison of the two architectures, helping determine whether SSMs can serve as viable building blocks for language modeling or require fundamental modifications.
The MMLU benchmark (Hendrycks et al., 2021) provides a particularly sharp lens into this performance gap. While SSMs can match Transformer results on many knowledge benchmarks, they tend to perform worse on MMLU unless trained significantly longer (Waleffe et al., 2024). This gap may stem from MMLU’s evaluation format: rather than scoring the likelihood of full-text answer candidates, models are required to select a letter label (A, B, C, or D) corresponding to the correct option. This format introduces an additional layer of complexity—models must not only possess the knowledge but also retrieve the corresponding letter from the context, creating a dual challenge of both knowledge and retrieval (Lieberum et al., 2023). To separate these distinct challenges, we analyze how each architecture handles the task. We find a clear division of roles: while factual knowledge is extracted from the model’s weights in a distributed manner across layers and heads, both Transformer- and SSM-based models rely on a small number of heads to retrieve the correct label from the input context.
@@ 1 Introduction
This sharp separation between knowledge and in-context retrieval underscores that it is not knowledge alone, but the skill of retrieving it from context, that drives success on benchmarks where SSMs typically fall short. Disabling retrieval heads leads to steep performance drops across architectures on tasks that require accessing and reusing prior information—such as mathematical reasoning (GSM8K), multi-step logical reasoning (BBH), and dialogue formats (ARC-C Chat)—while other benchmarks remain largely unaffected. This effect is particularly evident in MMLU, where in-context retrieval capabilities appear to be more important than stored knowledge. For example, our experiments with a pruned 4B-parameter Llama-3.1-8B submodel show that despite performing poorly on most benchmarks, it still achieves a strong MMLU score of 66%. However, when we disable just a single retrieval head, MMLU accuracy plummets to 25%, approximately the level of random guessing. This dramatic decline reveals that MMLU success depends more on retrieval capabilities—where Transformers maintain a clear advantage over SSMs—than on stored knowledge.
@@
The critical impact of in-context retrieval motivates our investigation of the underlying mechanism of these heads. We find that both Transformer- and SSM-based language models converge on a similar retrieval framework: a Gather-and-Aggregate (G&A) mechanism. Specifically, a Gather Head identifies and condenses segments of input tokens into a single vector representation (Figure 1(a)), which an Aggregate Head subsequently processes to extract the necessary information (Figure 1(b)). This finding indicates that, despite architectural differences, the two language models rely on similar retrieval strategies rather than fundamentally different approaches. Consequently, their retrieval gap reduces to the effectiveness with which a few specific heads implement the G&A mechanism.
@@
Building on this observation, we compare the G&A implementation of the two language models. As expected, we find that the known retrieval challenges of SSMs manifest in language models through the way a few key heads implement G&A. SSMs exhibit smoother attention patterns, making it difficult to execute the sharp token shifts needed for G&A, requiring more of their instances to achieve comparable performance. Moreover, we show how hybrid models (Glorioso et al., 2024; Lieber et al., 2024), which integrate a few attention layers into predominantly SSM-based architectures, overcome SSM’s limitations. When trained from scratch, hybrid models naturally delegate Aggregate Head functionality to attention layers. Conversely, in pretrained SSM models, replacing a single G&A instance with an attention-based one yields a significantly greater MMLU improvement than any other attention placement. While these results suggest that attention has a higher capacity for implementing G&A, they simultaneously highlight the mechanism’s concentrated nature. This indicates that not every layer needs to cache the full history—paving the way for more efficient architectures that store past tokens in fewer components.
@@
    Gather-and-Aggregate (G&A) Retrieval Framework. We formalize a unified mechanism—comprising Gather Heads and Aggregate Heads—that enables in-context retrieval, extending previous work by Lieberum et al. (2023); Wu et al. (2024) from Transformers to SSM-based language models and advancing mechanistic interpretability of SSMs. Our analysis reveals that G&A functionality is highly concentrated in SSMs as well, with just a few heads among thousands performing this function. Furthermore, we demonstrate that SSMs’ inherent retrieval limitations significantly impair their ability to implement G&A as effectively as Transformers.
@@
    Empirical Analysis of G&A across Benchmarks. We conduct extensive empirical analysis across Transformers, SSMs, and hybrid architectures, revealing how G&A heads critically determine performance on retrieval-intensive benchmarks such as MMLU, GSM8K, and BBH. Our findings suggest that the broader performance gap between Transformers and SSMs may stem from specific algorithmic limitations in a small subset of heads rather than general language modeling deficiencies. Most notably, we show that the difficulty of MMLU arises more from in-context retrieval demands than from knowledge representation, as models can perform well on MMLU despite struggling with other tasks, but fail catastrophically without functional G&A mechanism.
@@
    Understanding the Effectiveness of Hybrid Models: We demonstrate that hybrid models address the retrieval limitations of SSMs in implementing G&A. When trained from scratch, these models naturally assign Aggregate Head functionality to attention layers. Alternatively, “hybridizing” a pretrained SSM by replacing a single G&A component with an attention-based G&A significantly improves performance. Furthermore, we provide evidence that both architectures tend to develop these heads in the middle of the network, suggesting an effective placement for such hybrids.
@@ 2 Related Work
Sequence models, particularly Transformer-based language models with self-attention (Vaswani et al., 2023; Brown et al., 2020; Touvron et al., 2023), dominate modern language modeling due to their scalability and strong performance. Recently, recurrent models, like structured state-space models (SSMs) (Gu et al., 2022; Gu and Dao, 2023; Dao and Gu, 2024), have gained focus for efficiently processing long sequences. These models replace quadratic attention with sub-quadratic methods, offering competitive performance while maintaining efficiency. Other recurrent models with similar characteristics include (Dao et al., 2022; Sun et al., 2023; Katharopoulos et al., 2020; Yang et al., 2024a; Zhang et al., 2024b; Beck et al., 2024; Qin et al., 2024; Peng et al., 2024; Yang et al., 2024b). In this work, we use “SSM” to refer to these models as well, as our findings are broadly applicable.
@@
Throughout this study, we examine three different architectures: Llama-3.1-8B-Instruct (Touvron et al., 2023), a standard Transformer that achieves high accuracy on various NLP benchmarks; Falcon-Mamba-7B-Instruct (Zuo et al., 2024), a Mamba-1 SSM model without MLP layers; and Llamba-8B-Untied (Bick et al., 2025), a Mamba-2 SSM model distilled from Llama-3.1-8B using MOHAWK (Bick et al., 2024), which alternates MLP layers with SSMs. We refer to these as Llama-3.1-8B, Llamba-8B, and Falcon-Mamba-7B, respectively. Furthermore, Section 5 examines Zamba-2-7B (Glorioso et al., 2024), a hybrid model where Mamba2 layers outnumber attention layers by a 6:1 ratio, with only two attention layers shared across the entire model.
Central to understanding the performance differences between these architectures is the concept of token mixing. Token mixing refers to how sequence models propagate and combine information across different token positions. In self-attention, each token selectively incorporates information from other tokens based on learned weights. SSMs, in contrast, mix tokens using structured recurrences that enforce a fixed computation pattern. Ali et al. (2024); Dao and Gu (2024) link SSM’s token-mixing to the matrix-multiplication form of self-attention. They showed that Mamba-1 mimics thousands of attention heads with a head dimension of 1, interacting with the input in a multi-value manner. Dao and Gu (2024) introduced Mamba-2, which uses fewer heads with larger dimensions (Dao and Gu, 2024; Bick et al., 2025; Waleffe et al., 2024), making each head more expressive than in Mamba-1. These insights enable us to view all mentioned sequence models as structured matrices operating along the time dimension. For that reason, we generalize token-mixing mechanisms beyond traditional self-attention, and use throughout this work the term Temporal Mixing Heads (or simply Mixing Heads).
@@
SSM-based language models rely on compressed state representations, where information is continually updated in a low-dimensional hidden state—unlike Transformers, which cache their full context. These inherent differences have been observed across multiple benchmarks. Waleffe et al. (2024) found that Mamba-based models require extensive training to perform well on MMLU, while Wen et al. (2024) identified weaknesses in in-context retrieval tasks, such as associative recall. Jelassi et al. (2024) theoretically and empirically confirmed that SSMs have difficulty with precise copying. Arora et al. (2023) further demonstrated this gap, finding that associative recall capabilities account for majority of the performance difference between attention-based and gated-convolution models. Park et al. (2024) benchmarked Mamba and Transformer models across a range of in-context learning tasks, showing that SSMs underperform on tasks requiring non-standard retrieval, and proposed a hybrid architecture to address this limitation. Arora et al. (2025) analyzed the recall-throughput tradeoff across architectures, showing that attention-based models occupy a different region of the tradeoff frontier than SSMs, with stronger recall but higher computational cost. Blouir et al. (2025) complemented this view by enhancing retrieval in SSMs through training improvements without architectural changes. Unlike previous work, our study reveals that while these limitations exist, they originate from a small subset of heads responsible for retrieval rather than the entire language model.
@@
Most interpretability studies focus on attention heads in LLMs (Elhage et al., 2021; Todd et al., 2024; Lieberum et al., 2023; Zhang et al., 2024a; Olsson et al., 2022; Yu et al., 2024; Merullo et al., 2024; Rai et al., 2024; Zheng et al., 2024; Wu et al., 2024; Tulchinskii et al., 2024; Sanford et al., 2024), while relatively little attention has been given to recurrent mechanisms (Ali et al., 2024). Olsson et al. (2022) and Elhage et al. (2021) showed that induction heads emerge to support in-context learning. Subsequent work traced their training dynamics (Bietti et al., 2023), and mapped factual recall circuits (Meng et al., 2023).
In a more focused study, Lieberum et al. (2023) analyzed Chinchilla 70B on multiple-choice tasks and identified two key attention heads: the Content Gatherer Head and the Correct Letter Head, which work in conjunction with MLPs to select correct labels. In addition, Wu et al. (2024) identified Retrieval Heads—a small subset of Transformer heads essential for retrieval that comprise less than 5% of all heads but whose removal significantly disrupts retrieval, chain-of-thought reasoning, and increases hallucinations.
@@
Our contribution refines the understanding of their roles in supporting retrieval. We show that in both Transformers and SSMs, retrieval emerges not from a single head but through a coordinated mechanism: one head gathers the location of the target, and another aggregates the content. This framing extends prior work in three key ways. First, the “Correct Letter” heads identified by Lieberum et al. (2023) exhibit broader retrieval behavior than previously reported, going beyond multiple-choice selection. Second, retrieval depends on coordinated interactions between heads, challenging the interpretation in Wu et al. (2024) that treated individual heads as independently responsible. Third, we demonstrate that this mechanism also arises in SSMs—despite their lack of attention—highlighting its architectural generality. We introduce the term Aggregate Head to better reflect these empirical findings, while retaining “Gather” as shorthand for the “Content-Gatherer” heads described Lieberum et al. (2023).
@@
Concurrent with our work, Wiegreffe et al. (2025) expanded on Lieberum et al. (2023)’s findings beyond MMLU and Chinchilla-70B, investigating model behavior throughout training and analyzing cases of poor performance. They created synthetic tasks to separate knowledge from multiple-choice formatting abilities and uncovered mechanisms for how models adapt to different answer choice symbols. While their objectives align with ours, our research specifically targets the “algorithmic” in-context retrieval aspect of multiple-choice formatting across various architectures, contrasting this with models’ underlying knowledge capabilities.

### v2

@@ Abstract
State-space models (SSMs) offer efficient alternatives to Transformers for long sequences, but their fixed-size recurrent state limits capability on algorithmic tasks, such as retrieving past context. In this work, we examine how in-context retrieval operates in Transformer- and SSM-based language models and find that both rely on a similar Gather-and-Aggregate (G&A) mechanism: a Gather Head extracts relevant information pieces from context, which an Aggregate Head integrates into a single representation. In both architectures, G&A concentrates in a few heads, forming critical bottlenecks even for simple retrieval. For example, we show that disabling a single Gather or Aggregate Head in a pruned Llama-3.1-8B impairs retrieving the correct answer letter in MMLU, reducing its accuracy from 66% to 25% (random guessing). Moreover, this retrieval bottleneck can obscure limited knowledge demands of tasks as the pruned model succeeds on MMLU with functioning G&A heads yet fails on other knowledge benchmarks. The bottleneck similarly extends to tasks where SSMs typically underperform, such as GSM8K, BBH, and dialogue comprehension. We show that SSMs’ retrieval challenges manifest in these heads, creating smoother attention patterns instead of the sharp token transitions effective G&A requires. Thus, the Transformer-SSM retrieval gap exists in just a few heads, rather than the entire language model. This suggests a unified explanation for Transformer vs. SSM performance gap while showing how to merge their strengths. We find that pretrained hybrid models, where SSMs are combined with a few attention layers, delegate the role of Aggregate Heads to attention. Similarly, replacing a single G&A head in a pretrained SSM with an attention variant boosts retrieval and benchmark scores. Experiments are publicly available. [^1] https://github.com/goombalab/Gather-and-Aggregate
@@
1 Introduction
Transformers have driven major breakthroughs in language modeling, but their quadratic scaling with sequence length has renewed interest in recurrent alternatives like state-space models (SSMs). SSMs offer linear scaling with a fixed-size memory while remaining competitive across many tasks. However, replacing a large, explicit history cache with a compact state representation poses a fundamental tradeoff. SSM-based language models often struggle with the skill of “in-context retrieval”—the ability to precisely locate and extract tokens presented earlier in the sequence, such as recalling a phone number mentioned several paragraphs ago (Jelassi et al., 2024). This “algorithmic” challenge, which Transformers handle effectively with direct access to past tokens, has been identified as a key factor in the performance gap between Transformer models and SSM variants (Arora et al., 2023).
Previous work on SSM retrieval limitations has treated language models as black boxes or focused on simplified theoretical setups (Wen et al., 2024; Jelassi et al., 2024). Meanwhile, Transformer research shows that algorithmic capabilities, including retrieval, often concentrate in specific attention heads (Lieberum et al., 2023; Wu et al., 2024). This raises a key question: Can the Transformer-SSM performance gap be traced to algorithmic capabilities in a few specific heads? Clarifying this will enable a more targeted comparison of the two architectures, helping determine whether SSMs can serve as viable building blocks for language modeling or require fundamental modifications.
The MMLU benchmark (Hendrycks et al., 2021) exemplifies this performance gap. While SSMs match Transformers on many knowledge benchmarks, they significantly underperform on MMLU unless trained substantially longer (Waleffe et al., 2024). This gap stems from MMLU’s evaluation format: rather than scoring the likelihood of full-text answers, models must select letter labels (A, B, C, or D) corresponding to the correct option. This creates a dual challenge requiring both factual knowledge and retrieval capabilities (Lieberum et al., 2023).
To isolate these distinct challenges, we show that both architectures have a clear division of labor: factual knowledge extracted across layers and heads, but only a small number of specialized heads retrieve the correct labels from input context. This becomes particularly evident in our experiments where a 4B-parameter pruned Llama-3.1-8B achieves a strong MMLU score of 66%, which plummets to 25% when disabling a single retrieval head. Moreover, these heads obscure the true knowledge demands of such tasks—despite strong MMLU performance, the pruned model performs poorly on knowledge-focused benchmarks. Similar dependence extends beyond MMLU to mathematical reasoning (GSM8K), multi-step logical reasoning (BBH), and dialogue formats (ARC-C Chat), demonstrating that retrieval capabilities, not knowledge deficits, drive SSM underperformance in these domains.
@@ 1 Introduction
To understand how these specialized heads function, we investigate their underlying mechanism. We find that both Transformer- and SSM-based language models converge on a similar retrieval framework: a Gather-and-Aggregate (G&A) mechanism. Specifically, a Gather Head identifies and condenses segments of input tokens into a single vector representation (Figure 1(a)), which an Aggregate Head subsequently processes to extract the necessary information (Figure 1(b)). This finding indicates that the two language models rely on similar retrieval strategies, despite architectural differences. Consequently, their retrieval gap reduces to the effectiveness with which a few specific heads implement the G&A mechanism.
@@
Comparing G&A implementations, we find SSMs’ known retrieval challenges manifest in how key heads implement G&A. SSMs exhibit smoother attention patterns, making sharp token shifts difficult and requiring more heads to achieve comparable performance. Moreover, we show how hybrid models (Glorioso et al., 2024; Lieber et al., 2024), which integrate a few attention layers into predominantly SSM-based architectures, overcome SSM’s limitations. When trained from scratch, hybrid models naturally delegate Aggregate Head functionality to attention layers. Conversely, in pretrained SSM models, replacing a single G&A instance with an attention-based one yields a significantly greater MMLU improvement than any other attention placement. While attention appears better suited for G&A, the results also highlight the mechanism’s concentrated nature, suggesting that efficient architectures can store past tokens in fewer components.
@@
    Gather-and-Aggregate (G&A) Retrieval Framework. We formalize a unified mechanism—comprising Gather Heads and Aggregate Heads—that enables in-context retrieval, extending previous work by Lieberum et al. (2023); Wu et al. (2024) from Transformers to SSMs. Our analysis reveals that G&A functionality is highly concentrated in SSMs as well, with just a few heads among thousands performing this function. Furthermore, we demonstrate that SSMs’ inherent retrieval limitations significantly impair their ability to implement G&A as effectively as Transformers.
@@
    Empirical Analysis of G&A across Benchmarks. We analyze Transformers, SSMs, and hybrid architectures, showing that a small set of G&A heads critically drives performance on retrieval-heavy benchmarks such as GSM8K, BBH, and dialogue comprehension. This pattern is especially evident in MMLU, where we show that the difficulty stems more from in-context retrieval than from knowledge: models can succeed on MMLU despite struggling elsewhere, yet fail completely without functional G&A heads. These results suggest the Transformer–SSM gap stems from algorithmic limits in a few specific heads, not broad modeling deficiencies.
@@
    Understanding the Effectiveness of Hybrid Models: Hybrid models address the G&A limitations of SSMs by offloading Aggregate Head functionality to attention layers. When trained from scratch, attention heads often emerge near the middle of the network, suggesting a natural fit for aggregation. In pretrained SSMs, replacing a single G&A head with attention significantly improves performance, highlighting effective placement for such hybrids.
@@ 2 Related Work
Transformer-based models with self-attention (Vaswani et al., 2023; Brown et al., 2020; Touvron et al., 2023) dominate language modeling for their scalability and performance. Recently, recurrent models like structured state-space models (SSMs) (Gu et al., 2022; Gu and Dao, 2023; Dao and Gu, 2024) gained interest for their sub-quadratic efficiency and competitive results on long sequences. Other recurrent models include (Dao et al., 2022; Sun et al., 2023; Katharopoulos et al., 2020; Yang et al., 2024a; Zhang et al., 2024b; Beck et al., 2024; Qin et al., 2024; Peng et al., 2024; Yang et al., 2024b). We use ”SSM” to refer to these models as well, since our findings broadly apply to them.
@@
Throughout this study, we examine four architectures: Llama-3.1-8B-Instruct (Touvron et al., 2023), a standard Transformer; Falcon-Mamba-7B-Instruct (Zuo et al., 2024), a Mamba-1 SSM without MLP layers; Llamba-8B-Untied (Bick et al., 2025), a Mamba-2 SSM distilled from Llama-3.1-8B using MOHAWK (Bick et al., 2024); and Zamba-2-7B-Instruct (Glorioso et al., 2024), a hybrid model with a 6:1 ratio of Mamba2 to attention layers and only two shared attention layers. For brevity, we refer to these as Llama-3.1-8B, Falcon-Mamba-7B, Llamba-8B, and Zamba-2-7B, respectively.
Central to understanding the performance differences between these architectures is token mixing: how models propagate and combine information across token positions. In self-attention, tokens selectively aggregate context via learned weights, while SSMs use structured recurrences with more restricted patterns. Ali et al. (2024); Dao and Gu (2024) connect SSM token mixing to attention’s matrix-multiplication form, showing Mamba-1 mimics many low-dimensional attention heads. Mamba-2 (Dao and Gu, 2024; Bick et al., 2025; Waleffe et al., 2024) uses fewer, larger heads, increasing expressivity. These insights allow us to view all sequence models as structured matrices over time. Thus, we generalize token mixing beyond self-attention, referring to these mechanisms as Temporal Mixing Heads (or simply Mixing Heads).
@@
SSMs rely on compressed, continuously updated hidden states, unlike Transformers, which cache full context. This structural difference impacts retrieval performance across benchmarks. Waleffe et al. (2024) found that Mamba-based models require extensive training to perform well on MMLU, while Wen et al. (2024) identified weaknesses in in-context retrieval tasks, such as associative recall. Jelassi et al. (2024) theoretically and empirically confirmed that SSMs have difficulty with precise copying. Arora et al. (2023) demonstrated that associative recall capabilities account for the majority of performance differences between attention-based and gated-convolution models. Park et al. (2024) shows that SSMs underperform on tasks requiring non-standard retrieval, and proposed a hybrid architecture to address this limitation. Arora et al. (2025b) highlighted a recall-throughput tradeoff, with attention offering stronger recall at higher cost. Blouir et al. (2025) complemented this view by enhancing retrieval in SSMs through training improvements without architectural changes. Unlike previous work, we show these limitations originate from a small subset of retrieval heads rather than the entire model.
@@
Most interpretability studies focus on attention heads in LLMs (Elhage et al., 2021; Lieberum et al., 2023; Zhang et al., 2024a; Olsson et al., 2022; Yu et al., 2024; Merullo et al., 2024; Rai et al., 2024; Zheng et al., 2024; Wu et al., 2024; Tulchinskii et al., 2024; Sanford et al., 2024), while relatively little attention has been given to recurrent mechanisms (Ali et al., 2024). Olsson et al. (2022) and Elhage et al. (2021) showed that induction heads emerge to support in-context learning. Subsequent work traced their training dynamics (Bietti et al., 2023), and mapped factual recall circuits (Meng et al., 2023).
Lieberum et al. (2023) identified two key heads in Chinchilla-70B—Content Gatherer and Correct Letter—that, with MLPs, support label prediction in multiple-choice tasks. Wu et al. (2024) found that fewer than 5% of heads, termed Retrieval Heads, are critical for retrieval, Chain-of-thought, and reasoning.
@@
We refine the understanding of how heads support retrieval by showing that, in both Transformers and SSMs, retrieval arises from a coordinated mechanism—not a single head. One head identifies the target location, while another aggregates its content. This view extends prior work in three ways. First, the “Correct Letter” heads identified by Lieberum et al. (2023) exhibit broader retrieval behavior than previously reported, going beyond multiple-choice selection. Second, retrieval depends on coordinated interactions between heads, challenging the interpretation in Wu et al. (2024) that treated individual heads as independently responsible. Third, we demonstrate that this mechanism also arises in SSMs—despite their lack of attention—highlighting its architectural generality. We introduce the term Aggregate Head to better reflect these empirical findings, while retaining “Gather” as shorthand for the “Content-Gatherer” heads described Lieberum et al. (2023).
@@
Concurrent with our work, two lines of research investigate similar mechanisms. Wiegreffe et al. (2025) extend Lieberum et al. (2023) by analyzing training dynamics and failure modes in multiple-choice tasks. They introduce synthetic setups to separate knowledge from answer formatting and study how models adapt to symbolic answer spaces. While their goals align with ours, they focus on symbolic reasoning and performance variation, whereas we study the algorithmic basis of in-context retrieval across architectures.
Arora et al. (2025a) demonstrate that Transformers achieve in-context retrieval through key–value associations, whereas SSMs rely on direct retrieval. However, we find that both Transformer- and SSM-based models develop the same G&A mechanism, rather than differing mechanistically. This apparent contradiction stems from a key difference in setup: they train and evaluate shallow models (up to three layers) on synthetic tasks, while we study pretrained language models on natural benchmarks—leading to different mechanistic conclusions. While this distinction highlights that synthetic setups may reveal mechanisms that do not generalize to broader language modeling, we believe they remain valuable for probing model expressivity, such as understanding G&A head structure across architectures.